# Bar formation and evolution in the cosmological context: Inputs from the Auriga simulations

### v1

@@ 1 Introduction
While the fraction of barred galaxies is high in the local Universe, it is generally found to decrease towards higher redshifts (Sheth et al. 2008; Melvin et al. 2014; Simmons et al. 2014, but see also Jogee et al. 2004). Studies using data from the Hubble Space Telescope find that bars are present in about $\sim 10\%$ of disc galaxies up to $z\sim 2$ (e.g. Simmons et al. 2014), while more recent studies using data from the James Webb Space Telescope (JWST) find a higher fraction of bars at these redshifts: $\sim$20at $z\sim 1-2$ and $\sim$15between $z\sim 2-3$ (Le Conte et al., 2024). Bars have also recently been identified out to redshifts of $z\sim 4$ (Costantin et al., 2023; Smail et al., 2023; Tsukui et al., 2024; Amvrosiadis et al., 2024). At the same time, studies exploring stellar populations in local galaxies find evidence of old bars in the local Universe (Gadotti et al. 2015; de Sá-Freitas et al. 2023; de Sà-Freitas et al. in prep.), suggesting that these structures are robust and long-lived. The decreasing fraction of bars at higher redshifts, hints at that the tight link between bar formation and the evolution of spiral galaxies, while their presence in the early Universe suggests that spiral galaxies might be more dynamically mature at these early times than previously thought (see also e.g. Rizzo et al. 2020; Lelli et al. 2021).
@@ 5.2 Bars at high-$z$
A number of observational studies have explored the fraction of barred galaxies between redshifts of $z=0-1$, with most studies finding a declining bar fraction as a function of redshift (Abraham et al., 1999; Sheth et al., 2008; Melvin et al., 2014; Simmons et al., 2014), although some studies find constant bar fractions (specifically for strong bars) (Jogee et al., 2004). Studies using data from the Hubble Space Telescope (HST) have found that by $z=1$, the bar fraction drops to $\sim 10\%$ (e.g. Simmons et al. 2014). A more recent study by Le Conte et al. (2024) – covering the redshift range $z=1-3$ with data from the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al. 2023) of the James Webb Space Telescope – found that compared to previous HST results, the fraction of bars is higher by a factor of $\sim 2$ when considering the higher sensitivity JWST data, finding a bar fraction of $20\%$ at $z\sim 1-2$. At $z\sim 2-3$ the bar fraction drops to $15\%$. As shown in Fig. 4, these trends are reproduced in the Auriga simulations, with a decreasing fraction of barred galaxies as a function of redshift. It is important to note that our sample covers a small mass range at $z=0$, and therefore our results do not make predictions for the behaviour of lower mass galaxies.
@@ Acknowledgements
FF thanks Marie Martig, Simon D. White, Federico Lelli and Kyle Oman for useful discussions. The authors thank David Campbell and Adrian Jenkins for generating the initial conditions and selecting the sample of the Auriga galaxies. FF is supported by a UKRI Future Leaders Fellowship (grant no. MR/X033740/1). RG acknowledges support from an STFC Ernest Rutherford Fellowship (ST/W003643/1). FAG acknowledges support from ANID FONDECYT Regular 1211370 and by the ANID BASAL project FB210003. FAG, acknowledges funding from the Max Planck Society through a “Partner Group” grant. FM acknowledges support from the European Union - NextGenerationEU, in the framework of the HPC project – “National Centre for HPC, Big Data and Quantum Computing” (PNRR - M4C2 - I1.4 - CN00000013 – CUP J33C22001170001).

### v2

@@ 1 Introduction
While the fraction of barred galaxies is high in the local Universe, it is generally found to decrease towards higher redshifts (Sheth et al. 2008; Melvin et al. 2014; Simmons et al. 2014, but see also Jogee et al. 2004). Studies using data from the Hubble Space Telescope find that bars are present in about $\sim 10\%$ of disc galaxies up to $z\sim 2$ (e.g. Simmons et al. 2014), while more recent studies using data from the James Webb Space Telescope (JWST) find multiple bars at $z>1$ (e.g. Guo et al. 2023; Costantin et al. 2023) and a higher bar fraction than previously reported: $\sim$20at $z\sim 1-2$ and $\sim$15between $z\sim 2-3$ (Le Conte et al., 2024). Bars have also recently been identified out to redshifts of $z\sim 4$ (Smail et al., 2023; Tsukui et al., 2024; Amvrosiadis et al., 2024). At the same time, studies exploring stellar populations in local galaxies find evidence of old bars in the local Universe (Gadotti et al. 2015; de Sá-Freitas et al. 2023; de Sà-Freitas et al. in prep.), suggesting that these structures are robust and long-lived. The decreasing fraction of bars at higher redshifts, hints at that the tight link between bar formation and the evolution of spiral galaxies, while their presence in the early Universe suggests that spiral galaxies might be more dynamically mature at these early times than previously thought (see also e.g. Rizzo et al. 2020; Lelli et al. 2021).
@@ 5.2 Bars at high-$z$
A number of observational studies have explored the fraction of barred galaxies between redshifts of $z=0-1$, with most studies finding a declining bar fraction as a function of redshift (Abraham et al., 1999; Sheth et al., 2008; Melvin et al., 2014; Simmons et al., 2014), although some studies find constant bar fractions (specifically for strong bars) (Jogee et al., 2004). Studies using data from the Hubble Space Telescope (HST) have found that by $z=1$, the bar fraction drops to $\sim 10\%$ (e.g. Simmons et al. 2014). Recent studies using JWST data have been revealing barred galaxies at $z>1$ (e.g. Guo et al. 2023), and shedding new light on the bar fraction at these redshifts. For example, a recent study by Le Conte et al. (2024) – covering the redshift range $z=1-3$ with data from the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al. 2023) – found that compared to previous HST results, the fraction of bars is higher by a factor of $\sim 2$ when considering the higher sensitivity JWST data, finding a bar fraction of $20\%$ at $z\sim 1-2$. At $z\sim 2-3$ the bar fraction drops to $15\%$. As shown in Fig. 4, these trends are reproduced in the Auriga simulations, with a decreasing fraction of barred galaxies as a function of redshift. It is important to note that our sample covers a small mass range at $z=0$, and therefore our results do not make predictions for the behaviour of lower mass galaxies.
@@ Acknowledgements
The authors thank Marie Martig, Simon White, Federico Lelli and Kyle Oman for useful discussions, and David Campbell and Adrian Jenkins for generating the initial conditions and selecting the sample of the Auriga galaxies. FF is supported by a UKRI Future Leaders Fellowship (grant no. MR/X033740/1). RG acknowledges support from an STFC Ernest Rutherford Fellowship (ST/W003643/1). FAG acknowledges support from ANID FONDECYT Regular 1211370, by the ANID BASAL project FB210003 and acknowledges funding from the Max Planck Society through a “Partner Group” grant. FM acknowledges support from the European Union - NextGenerationEU, in the framework of the HPC project – “National Centre for HPC, Big Data and Quantum Computing” (PNRR - M4C2 - I1.4 - CN00000013 – CUP J33C22001170001). This work was supported by STFC [ST/X001075/1]. Part of the simulations of this paper used: the SuperMUC system at the Leibniz Computing Centre (www.lrz.de), Garching, under the project PR85JE of the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu); the Freya computer cluster at the Max Planck Institute for Astrophysics; and the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.